If the price of indomie falls from 017 by 0.01 naira causing an increase in quantity demanded from 100 units to 110 units. Calculate the price elasticity of demand for indomie and interpret the result.


1
Expert's answer
2021-02-28T11:37:59-0500

By the definition of the price elasticity of demand, we have:


"E_d=\dfrac{\%\Delta Q}{\%\Delta P}=\dfrac{\dfrac{Q_2-Q_1}{Q_1}\times100\%}{\dfrac{P_2-P_1}{P_1}\times100\%},""E_d=\dfrac{\dfrac{110-100}{100}\times100\%}{\dfrac{0.16-0.17}{0.17}\times100\%}=\dfrac{10\%}{-5.88\%}=-1.70."

Since, "E_d\gt 1" (we do not bother about sign minus, we look at the coefficient of the price elasticity of demand), the demand is elastic.